OpenStudy (anonymous):

Help!! Solve for x: 5x + 2 = 4x - 9. 7 -7 11 -11

OpenStudy (solomonzelman):

5x + 2 = 4x - 9 5x + 2 - 4x = 4x - 9 -4x x + 2 = -9 x + 2 -2 = -9 - 2 x = -11 CHECK, 5x + 2 = 4x - 9 5(-11) + 2 = 4(-11) - 9 -55+2=-44-9 -53=-53 Yes, x= -11

OpenStudy (anonymous):

thanks